# SPAGS: Sparse-View Articulated Object Reconstruction from Single State via Planar Gaussian Splatting

### v2

@@ 3.3 Planar Gaussian Refinement via Few-shot Diffusion
To augment the training data, we employ the leave-one-out strategy to generate additional image pairs, similar as [yang2024gaussianobject]. We divide the input data into $n$ subsets, each of which contains n1n-1 original images and one left-out image $x_{\text{left}}$. For each subset, we train the planar Gaussians for 6,000 iterations without $x_{\text{left}}$ to yield the degraded renderings $x_{\text{deg}}$ and $\bar{x}_{\text{deg}}$ for $x_{\text{left}}$ and its corresponding generated label $\bar{x}_{\text{left}}$ respectively. Subsequently, we incorporate the left-out image $x_{\text{left}}$ to continue training the Gaussian primitives 4,000 iterations, producing the repaired rendering for $x_{\text{deg}}$ and $\bar{x}_{\text{deg}}$. These renderings at different iterations are combined with $x_{\text{left}}$ and $\bar{x}_{\text{left}}$ to form image pairs for diffusion model fine-tuning.
@@
Once the diffusion model fine-tuning is completed, we utilize it to optimize the noise regions $\Omega_{\text{noisy}}$. We start with randomly sampling viewpoints $\sigma_{\text{rand}}$ from $\Omega_{\text{noisy}}$, and obtain corresponding rendering $x(\mathcal{G},\sigma_{\text{rand}})$. Following [meng2022sdedit], we use a latent diffusion encoder $\boldsymbol{\varepsilon}$ to encode the rendering, later employ the image condition to yield a noisy latent representation $\mathbf{z}_{t}$:
@@
where $\Theta$ denotes the back-projection operation via the depth rendering, and $H$ is the homography matrix follwing [chen2024pgsr], mapping $\delta_{i}$ to its corresponding pixel batch under $\bar{\sigma}_{i}$. In this way, we enhance the geometric consistency of the planar Gaussians. The total training objective of the refinement is formulated as: $\mathcal{L}_{\text{refine}}=\mathcal{L}_{\text{repair}}+\lambda_{\text{vc}}\mathcal{L}_{\text{vc}}+\mathcal{L}_{\text{planar}}$.
3.4 Articulation Modeling for Planar Gaussians
After the coarse-to-fine optimization, we employ GPT-4o to generate an abstract articulated tree for planar Gaussians, taking multiple rendered images as prompts. The articulated tree describes the name and connectivity of each part as well as the joint type.
Subsequently, we introduce a part-aware probability $m_{i}$ for each Gaussian primitive $\mathcal{G}_{i}$. We render Gaussian primitives from a series of viewpoints to yield multiple rendered images, then we employ the Lang-SAM [lang_segment_anything] to generate 2D part-level segmentation masks $\mathcal{M}$ for these renderings, using the part names from the articulated tree as text prompts. Following [chen2023gaussianeditor], the probability $m_{i}^{o}$ of $\mathcal{G}_{i}$ belongs to part $o$ can be derived through back-projecting $\mathcal{M}$ of corresponding pixel $\rho$:
@@ 3.4 Articulation Modeling for Planar Gaussians
After obtaining the part-level Gaussian primitives, we first determine a connecting area between two neighboring parts through a distance threshold from Gaussian center to its neighboring part. Then we use the renderings of these clusters and the articulated tree as prompts to query VLM for joint parameters, as shown in Fig. 3. Please refer to the supplementary materials for a detailed elaboration.
Table 2: Quantitative results for the novel view synthesis on PartNet-Mobility dataset. denotes methods designed for multi-view reconstruction. We bold the best results and underline the second best results.
@@
PSNR $\uparrow$ PGSR [chen2024pgsr] 19.68 19.55 24.75 11.90 12.70 17.60 27.82 22.69 17.36 17.85 19.19
CoherentGS [paliwal2024coherentgs] 11.54 11.99 13.60 12.85 9.31 15.12 16.88 12.64 13.63 7.93 12.55
Sparse2DGS [Wu_2025_CVPR] 21.80 17.51 25.15 22.00 22.29 19.79 27.79 16.07 19.12 22.21 21.37
GaussianObject [yang2024gaussianobject] 23.42 22.17 25.52 23.11 20.94 19.04 28.85 21.46 20.53 14.41 21.95
@@
SSIM $\uparrow$ PGSR [chen2024pgsr] 0.93 0.91 0.95 0.63 0.76 0.89 0.97 0.94 0.84 0.86 0.87
CoherentGS [paliwal2024coherentgs] 0.85 0.83 0.92 0.82 0.76 0.86 0.94 0.91 0.77 0.57 0.82
Sparse2DGS [Wu_2025_CVPR] 0.96 0.90 0.96 0.83 0.89 0.70 0.98 0.91 0.89 0.91 0.89
GaussianObject [yang2024gaussianobject] 0.96 0.92 0.96 0.92 0.91 0.88 0.98 0.91 0.87 0.78 0.90
@@
LPIPS $\downarrow$ PGSR [chen2024pgsr] 0.07 0.10 0.06 0.43 0.22 0.13 0.02 0.07 0.19 0.22 0.15
CoherentGS [paliwal2024coherentgs] 0.17 0.19 0.10 0.26 0.25 0.20 0.09 0.09 0.27 0.53 0.22
Sparse2DGS [Wu_2025_CVPR] 0.06 0.09 0.12 0.12 0.10 0.11 0.03 0.14 0.12 0.15 0.10
GaussianObject [yang2024gaussianobject] 0.04 0.12 0.03 0.09 0.07 0.15 0.02 0.09 0.14 0.29 0.10
@@
Table 3: Quantitative results for part-level mesh reconstruction on PartNet-Mobility dataset. CD-d and CD-s represent the CD results for dynamic parts and static parts respectively.
CD-d $\downarrow$ REArtGS [wu2025reartgs] 98.57 56.28 32.41 128.09 204.30 473.19 131.76 63.37 489.21 171.98 184.92
ArtGS [liu2025building] 137.51 28.47 49.75 158.24 134.29 507.07 85.84 104.01 543.75 178.21 192.71
SPAGS (Ours) 27.74 6.00 6.26 23.94 0.95 5.69 17.85 8.44 95.14 66.72 25.87
CD-s $\downarrow$ REArtGS [wu2025reartgs] 89.51 52.20 107.36 85.47 103.98 184.71 26.44 218.51 115.60 36.69 102.05
ArtGS [liu2025building] 62.65 64.01 77.51 99.61 195.07 271.24 16.27 192.04 100.34 25.10 110.38
SPAGS (Ours) 66.82 42.01 8.80 48.63 23.44 18.25 1.80 21.75 53.74 24.92 31.02
4 Experiments
4.1 Experimental Setting
The comparisons include SOTA methods: PGSR [chen2024pgsr] (planar Gaussians), ArtGS [liu2025building], REArtGS [wu2025reartgs] (category-agnostic reconstruction), CoherentGS [paliwal2024coherentgs], Sparse2DGS [Wu_2025_CVPR], GaussianObject [yang2024gaussianobject] (sparse-view reconstruction). Except for ArtGS and REArtGS, all other methods are trained using the same 4-view RGB images as ours, while ArtGS and REArtGS are provided with additional 4-view images from another motion state.
@@ 4.1 Experimental Setting
Datasets. We conduct our experiments on both synthesis and real-world data. For synthesis data, we select the same 10 categories from PartNet-Mobility [xiang2020sapien] as REArtGS. Each object contains 64 randomly sampled candidate camera poses and corresponding RGB images. For real-world data, we scan 6 categories of articulated objects, characterized by a diverse scale and geometry. Each object contains 76 to 160 randomly sampled candidate camera poses and corresponding RGB images captured in unbounded indoor scenes. We apply the optimal view selection algorithm to determine 4 optimal sparse views for each object in the two datasets, and use SAM [kirillov2023segany] to generate object masks in real-world data. Note that we only use the candidate camera poses without images in view perception.
Table 4: Quantitative results for mesh reconstruction on real-world data.
Metrics Method Stapler USB Scissor Knife Drawer Bin Mean
CD $\downarrow$ CoherentGS [paliwal2024coherentgs] 30.65 10.92 13.12 8.76 104.02 64.84 38.72
Sparse2DGS [Wu_2025_CVPR] 34.07 7.63 15.24 3.21 84.15 65.38 34.95
GaussianObject [yang2024gaussianobject] 11.24 3.09 5.23 4.76 21.54 15.30 10.19
SPAGS (Ours) 3.13 2.34 5.76 2.82 16.36 4.00 5.74
F1 $\uparrow$ CoherentGS [paliwal2024coherentgs] 0.11 0.14 0.02 0.16 0.01 0.02 0.08
Sparse2DGS [Wu_2025_CVPR] 0.14 0.20 0.02 0.21 0.03 0.03 0.11
GaussianObject [yang2024gaussianobject] 0.16 0.25 0.05 0.22 0.01 0.07 0.13
SPAGS (Ours) 0.18 0.24 0.10 0.27 0.03 0.07 0.15
EMD $\downarrow$ CoherentGS [paliwal2024coherentgs] 0.25 0.23 0.35 0.15 0.64 0.59 0.37
Sparse2DGS [Wu_2025_CVPR] 0.16 0.18 0.23 0.18 0.58 0.50 0.31
GaussianObject [yang2024gaussianobject] 0.09 0.11 0.16 0.14 0.23 0.27 0.17
SPAGS (Ours) 0.06 0.07 0.10 0.07 0.13 0.08 0.09
@@
Figure 4: The qualitative results of whole mesh reconstruction on PartNet-Mobility dataset.
@@
Figure 5: The qualitative results of novel view synthesis on PartNet-Mobility dataset.
Figure 6: The qualitative results of articulated modeling. We set the dynamic parts to yellow and the static parts to blue, and use red arrows to represent the joints.
4.2 Comparison with State-of-the-Art Methods
@@
Mesh Reconstruction. We use the Chamfer distance (CD), F1-score, and EMD [zhang2022deepemd] as the metrics for surface mesh quality. We provide the quantitative results in Table. 1 and present qualitative results in Fig. 4. More qualitative results can be found in the supplementary materials. Our method achieve bests mean results across all metrics with 9.60, 0.17 and 0.12, outperforming other methods significantly. We observe that Sparse2DGS exhibits lots of artifacts under sparse view supervisions, as illustrated in Fig. 4. GaussianObject and CoherentGS produce surface meshes with lots of artifacts, primarily caused by the lack of geometric constraints. In contrast, our method maintains high-fidelity reconstruction by the coarse-to-fine optimization of planar Gaussians.
@@
Novel View Synthesis Performance. We employ PSNR, SSIM, and LPIPS as metrics for novel view synthesis. Table. 2 shows the quantitative results, and Fig. 5 presents qualitative comparisons. More qualitative results can be found in the supplementary materials. Our method achieves best results across all metrics with 24.13, 0.94 and 0.07. As shown in Fig. 5, our method exhibits more realistic rendering results, while other methods produce significant artifacts and noise in renderings, especially for Sparse2DGS and CoherentGS. This is mainly because our method enhances noise region reconstruction through few-shot diffusion.
@@
Articulated Modeling. We denote the CD results for dynamic and static part meshes as CD-d and CD-s respectively. The quantitative results of part-level mesh reconstruction are presented in Table. 3 and the qualitative results are shown in Fig. 6. We also provide quantitative results of joint estimation in the supplementary materials. Although ArtGS and REArtGS utilize two-stage observations, our method still achieve bests mean results on both CD-d and CD-s with 25.87, 31.02, significantly exceeding them. As shown in Fig. 6, our method yields high-fidelity part-level mesh reconstruction and accurate joint estimation. This demonstrates our part segmentation and joint estimation for planar Gaussians are simple but effective, achieving realistic results without two-stage observations.
@@
Table 5: Ablation of view perception on PartNet-Mobility dataset. Mean results are reported.
Settings CD $\downarrow$ CD-d $\downarrow$ CD-s $\downarrow$
@@
Optimal Perception 9.60 27.87 44.91
4.3 Ablation Study
Ablation on View Perception. We conduct the ablation study of view perception on PartNet-Mobility dataset. Quantitative results are shown in Table. 5. “Random Sampling” denotes selecting 4-view images randomly from candidates and we take the average results over 10 trials. “Pre-defined” refers to manually choosing promising 4-view images covering 360 degrees around the objects.

### v4

@@ III-C Planar Gaussian Refinement via Few-shot Diffusion
Figure 4: Illustration of data augmentation via few-shot image inputs.
To augment the training data, we employ the leave-one-out strategy to generate additional image pairs, similar as [40]. We divide the input data into $n$ subsets, each of which contains n1n-1 original images and one left-out image $x_{\text{left}}$. For each subset, we train the planar Gaussians for 6,000 iterations without $x_{\text{left}}$ to yield the degraded renderings $x_{\text{deg}}$ and $\bar{x}_{\text{deg}}$ for $x_{\text{left}}$ and its corresponding generated label $\bar{x}_{\text{left}}$ respectively. Subsequently, we incorporate the left-out image $x_{\text{left}}$ to continue training the Gaussian primitives 4,000 iterations, producing the repaired rendering for $x_{\text{deg}}$ and $\bar{x}_{\text{deg}}$. These renderings at different iterations are combined with $x_{\text{left}}$ and $\bar{x}_{\text{left}}$ to form image pairs for diffusion model fine-tuning, as shown in Fig. 4.
@@
Once the diffusion model fine-tuning is completed, we utilize it to optimize the noise regions $\Omega_{\text{noisy}}$. We start with randomly sampling viewpoints $\sigma_{\text{rand}}$ from $\Omega_{\text{noisy}}$, and obtain corresponding rendering $x(\mathcal{G},\sigma_{\text{rand}})$. Following [21], we use a latent diffusion encoder $\boldsymbol{\varepsilon}$ to encode the rendering, later employ the image condition to yield a noisy latent representation $\mathbf{z}_{t}$:
@@
where $\Theta$ denotes the back-projection operation via the depth rendering, and $H$ is the homography matrix follwing [1], mapping $\delta_{i}$ to its corresponding pixel batch under $\bar{\sigma}_{i}$. In this way, we enhance the geometric consistency of the planar Gaussians. The total training objective of the refinement is formulated as: $\mathcal{L}_{\text{refine}}=\mathcal{L}_{\text{repair}}+\lambda_{\text{vc}}\mathcal{L}_{\text{vc}}+\mathcal{L}_{\text{planar}}$.
TABLE I: Exemplary Visual Question Answering (VQA) Sessions with the VLM.
Task Objective VQA Interaction Examples
Identify the topology of
revolute joint
Prompt: Analyze the rendered image. Does the physical hinge connect two distinct components (e.g., a container and its lid), or is it a surface-mounted pivot (e.g., a knob)?
Response: The physical hinge establishes a connection between two distinct components. It is classified as type (1).
Revolute axis
candidate selection
(type 1)
Prompt: Observe the rendered grid image with numbered spatial anchors. Select $n_{2}\geq 2$ points from the clustered articulation region to define the primary rotation axis.
Response: Points 5 and 8 are spatially aligned with the physical hinge. I designate points [5, 8] to parameterize the rotation axis.
Revolute axis
candidate selection
(type 2)
Prompt: Based on the identified articulation region and its estimated planar normal vector, select the optimal anchor point to determine the pivot’s vertical axis.
Response: The geometric center of the rotating knob serves as the optimal pivot. I select point [12].
Identify the topology of
prismatic joint
Prompt: Analyze the rendering. Does this mechanism represent an orthogonal sliding joint (e.g., a drawer) or a tangential surface sliding joint (e.g., a sliding blade)?
Response: The component exhibits inward/outward translation relative to the base. It is identified as an orthogonal prismatic joint.
Surface sliding
direction alignment
Prompt: Evaluate the directional arrows superimposed on the rendered image. Identify the index of the arrow that accurately represents the physical sliding trajectory.
Response: The horizontal vector is congruent with the physical sliding track. The selected orientation corresponds to arrow [3].
III-D Articulation Modeling for Planar Gaussians
After the coarse-to-fine optimization, we employ GPT-4o to generate an abstract articulated tree for planar Gaussians, taking multiple rendered images as prompts, as shown in Fig. 5. The articulated tree describes the name and connectivity of each part as well as the joint type.
Subsequently, we introduce a part-aware probability $m_{i}$ for each Gaussian primitive $\mathcal{G}_{i}$. We render Gaussian primitives from a series of viewpoints to yield multiple rendered images, then we employ the Lang-SAM [20] to generate 2D part-level segmentation masks $\mathcal{M}$ for these renderings, using the part names from the articulated tree as text prompts. Following [2], the probability $m_{i}^{o}$ of $\mathcal{G}_{i}$ belongs to part $o$ can be derived through back-projecting $\mathcal{M}$ of corresponding pixel $\rho$:
@@ III-D Articulation Modeling for Planar Gaussians
TABLE II: Quantitative results for the surface reconstruction quality on PARIS dataset. denotes methods designed for multi-view reconstruction.
@@
CD $\downarrow$ PGSR [1] 13.19 26.94 6.01 672.76 32.20 32.40 2.49 24.12 48.68 81.90 94.07
CoherentGS [26] 324.65 33.76 313.04 40.87 80.61 64.31 9.34 164.08 41.42 303.67 137.58
Sparse2DGS [34] 49.25 35.57 28.96 38.87 12.64 425.27 4.13 28.47 73.74 60.88 75.78
GaussianObject [40] 7.39 6.92 0.82 36.05 0.58 68.87 1.58 166.03 12.30 126.93 42.75
SPAGS (Ours) 5.65 6.56 0.89 9.04 0.66 27.54 1.05 5.21 17.56 21.80 9.60
F1 $\uparrow$ PGSR [1] 0.08 0.13 0.33 0.02 0.15 0.05 0.25 0.19 0.03 0.01 0.12
CoherentGS [26] 0.01 0.06 0.08 0.06 0.15 0.03 0.32 0.17 0.02 0.00 0.09
Sparse2DGS [34] 0.11 0.07 0.24 0.05 0.18 0.03 0.16 0.13 0.03 0.04 0.10
GaussianObject [40] 0.10 0.15 0.38 0.05 0.35 0.05 0.27 0.10 0.03 0.01 0.15
SPAGS (Ours) 0.14 0.19 0.40 0.06 0.38 0.05 0.32 0.13 0.03 0.02 0.17
EMD $\downarrow$ PGSR [1] 0.14 0.19 0.07 0.84 0.16 0.17 0.06 0.13 0.22 0.23 0.22
CoherentGS [26] 0.53 0.26 0.44 0.20 0.32 0.23 0.06 0.16 0.25 0.60 0.31
Sparse2DGS [34] 0.50 0.59 0.64 0.62 0.30 0.72 0.50 0.51 0.28 0.31 0.50
GaussianObject [40] 0.17 0.12 0.09 0.12 0.09 0.20 0.05 0.31 0.17 0.38 0.17
SPAGS (Ours) 0.12 0.11 0.09 0.10 0.08 0.17 0.04 0.10 0.16 0.20 0.12
After obtaining the part-level Gaussian primitives, we utilize Visual Question Answering (VQA) to achieve open-vocabulary articulated parameter estimation, as shown in Table. I. Following Articulated-Anymesh [27], we categorize revolute joints via a VLM (GPT-4o [25]) into two types based on their topology: (1) hinge-connected parts and (2) surface-anchored axes (e.g., knobs). For type (1), we cluster $n_{1}\geq 10$ candidate points in the junction area, project labeled points onto 2D renderings, and prompt the VLM to select $n_{2}\geq 2$ points to define the axis, as shown in Fig. 6. The joint pivot is computed as the mean position, with the direction derived via Singular Value Decomposition (SVD) on the point offsets. For type (2), the joint is defined by the plane normal and a selected pivot point.
Similarly, prismatic joints are functionally classified into inward/outward or surface-sliding types. The former’s direction is determined by the fitted plane normal of the connection area, while the latter is identified through VLM-based selection among visual arrow prompts. Please refer to the supplementary materials for a detailed elaboration.
TABLE III: Quantitative results for the novel view synthesis on PARIS dataset. denotes methods designed for multi-view reconstruction.
PSNR $\uparrow$ PGSR [1] 19.68 19.55 24.75 11.90 12.70 17.60 27.82 22.69 17.36 17.85 19.19
CoherentGS [26] 11.54 11.99 13.60 12.85 9.31 15.12 16.88 12.64 13.63 7.93 12.55
Sparse2DGS [34] 21.80 17.51 25.15 22.00 22.29 19.79 27.79 16.07 19.12 22.21 21.37
GaussianObject [40] 23.42 22.17 25.52 23.11 20.94 19.04 28.85 21.46 20.53 14.41 21.95
@@
SSIM $\uparrow$ PGSR [1] 0.93 0.91 0.95 0.63 0.76 0.89 0.97 0.94 0.84 0.86 0.87
CoherentGS [26] 0.85 0.83 0.92 0.82 0.76 0.86 0.94 0.91 0.77 0.57 0.82
Sparse2DGS [34] 0.96 0.90 0.96 0.83 0.89 0.70 0.98 0.91 0.89 0.91 0.89
GaussianObject [40] 0.96 0.92 0.96 0.92 0.91 0.88 0.98 0.91 0.87 0.78 0.90
@@
LPIPS $\downarrow$ PGSR [1] 0.07 0.10 0.06 0.43 0.22 0.13 0.02 0.07 0.19 0.22 0.15
CoherentGS [26] 0.17 0.19 0.10 0.26 0.25 0.20 0.09 0.09 0.27 0.53 0.22
Sparse2DGS [34] 0.06 0.09 0.12 0.12 0.10 0.11 0.03 0.14 0.12 0.15 0.10
GaussianObject [40] 0.04 0.12 0.03 0.09 0.07 0.15 0.02 0.09 0.14 0.29 0.10
@@
TABLE IV: Quantitative results for articulated modeling on the PARIS dataset.
Metrics Method Stapler USB Scissors Fridge Foldchair Washer Oven Laptop Blade Storage Mean
Axis Ang. $\downarrow$ ArtGS [17] 30.14 76.48 84.35 44.51 18.63 78.05 86.37 15.48 26.81 62.35 52.32
REArtGS++ [32] 40.24 75.27 62.15 66.34 83.17 68.24 75.78 42.04 61.50 36.02 61.07
SPAGS (Ours) 4.01 0.07 0.03 0.02 0.25 5.13 7.48 0.04 3.11 0.14 2.03
Axis Pos. $\downarrow$ ArtGS [17] 0.16 0.31 0.24 0.54 0.09 0.16 0.39 0.07 - - 0.20
REArtGS++ [32] 0.19 0.25 0.27 0.59 0.11 0.24 0.43 0.05 - - 0.21
SPAGS (Ours) 0.09 0.01 0.05 0.02 0.03 0.11 0.02 0.01 - - 0.03
CD-s $\downarrow$ ArtGS [17] 62.65 64.01 77.51 99.61 195.07 271.24 100.34 192.04 16.27 25.10 110.38
REArtGS++ [32] 89.51 52.20 107.36 85.47 103.98 184.71 115.60 218.51 26.44 36.69 102.05
SPAGS (Ours) 66.82 42.01 8.80 48.63 23.44 18.25 53.74 21.75 1.80 24.92 31.02
CD-m $\downarrow$ ArtGS [17] 137.51 28.47 49.75 158.24 134.29 507.07 543.75 104.01 85.84 178.21 192.71
REArtGS++ [32] 98.57 56.28 32.41 128.09 204.30 473.19 489.21 63.37 131.76 171.98 184.92
SPAGS (Ours) 27.74 6.00 6.26 23.94 0.95 5.69 95.14 8.44 17.85 66.72 25.87
CD $\downarrow$ ArtGS [17] 85.12 58.60 23.45 120.30 45.70 201.10 185.00 110.20 48.90 130.40 100.90
REArtGS++ [32] 71.50 49.30 18.90 105.10 39.80 178.60 162.20 95.70 35.60 110.90 86.80
SPAGS (Ours) 5.65 6.56 0.89 9.04 0.66 27.54 17.56 5.21 1.05 21.80 9.60
IV Experiments
IV-A Experimental Setting
The comparisons include SOTA methods: PGSR [1] (planar Gaussians), ArtGS [17], REArtGS++ [32] (category-agnostic reconstruction), CoherentGS [26], Sparse2DGS [34], GaussianObject [40] (sparse-view reconstruction). Except for ArtGS and REArtGS++, all other methods are trained using the same 4-view RGB images as ours, while ArtGS and REArtGS++ are provided with additional 4-view images from another motion state.
For whole object reconstruction, we use Chamfer Distance (CD), F1-score, Earth Mover’s Distance (EMD) [42] as the metrics of mesh quality, and use PSNR, SSIM, LPIPS [44] as novel view synthesis metric. For articulated modeling, we employ CD-s and CD-m for surface meshes of static parts and dynamic parts respectively. We evaluate the predicted joints using angular error (Axis Ang.) and pivot distance (Axis Pos., for revolute joints only). Note that the CD resuls are multiplied by 1,000.
@@ IV-A Experimental Setting
Datasets. We conduct our experiments on both synthesis and real-world data. For synthesis data, we select the same 10 categories from PARIS dataset [35] and 5 categories from ArtGS-Multi dataset [17] as REArtGS++. Specifically, the PARIS dataset contains articulated objects composed of two parts, whereas the ArtGS-Multi dataset comprises multi-part articulated objects. Each object contains 64 randomly sampled candidate camera poses and corresponding RGB images.
@@
For real-world data, we scan 6 categories of articulated objects, characterized by a diverse scale and geometry. Each object contains 76 to 160 randomly sampled candidate camera poses and corresponding RGB images captured in unbounded indoor scenes. We apply the optimal view selection algorithm to determine 4 optimal sparse views for each object in the two datasets, and use SAM [12] to generate object masks in real-world data. Note that we only use the candidate camera poses without images in view perception.
@@
Figure 7: The qualitative results of whole mesh reconstruction on PARIS dataset.
@@
Figure 8: The qualitative results of novel view synthesis on PARIS dataset.
Figure 9: The qualitative results of articulated modeling on PARIS dataset.
IV-B Comparison with State-of-the-Art Methods
@@
Mesh Reconstruction. We use the Chamfer distance (CD), F1-score, and EMD [42] as the metrics for surface mesh quality. We provide the quantitative results in Table. II and present qualitative results in Fig. 7. More qualitative results can be found in the supplementary materials. Our method achieve bests mean results across all metrics with 9.60, 0.17 and 0.12, outperforming other methods significantly. We observe that Sparse2DGS exhibits lots of artifacts under sparse view supervisions, as illustrated in Fig. 7. GaussianObject and CoherentGS produce surface meshes with lots of artifacts, primarily caused by the lack of geometric constraints. In contrast, our method maintains high-fidelity reconstruction by the coarse-to-fine optimization of planar Gaussians.
@@
Novel View Synthesis Performance. We employ PSNR, SSIM, and LPIPS as metrics for novel view synthesis. Table. III shows the quantitative results, and Fig. 8 presents qualitative comparisons. More qualitative results can be found in the supplementary materials. Our method achieves best results across all metrics with 24.13, 0.94 and 0.07. As shown in Fig. 8, our method exhibits more realistic rendering results, while other methods produce significant artifacts and noise in renderings, especially for Sparse2DGS and CoherentGS. This is mainly because our method enhances noise region reconstruction through few-shot diffusion. Moreover, we provide the qualitative results of generation at unseen states, as shown in Fig. 10.
Figure 10: Additional qualitative results of generation at unseen states on PARIS dataset.
Figure 11: The qualitative results of articulated modeling on ArtGS-Multi dataset.
@@
Articulated Modeling. On the basis of the overall CD, we additionally compare CD-m, CD-s, Axis Ang. and Axis Pos., as illustrated in Sec. IV-A. The quantitative results of articulated modeling on PARIS dataset are presented in Table. IV and the qualitative results are shown in Fig. 9. We also provide quantitative results of joint estimation in the supplementary materials. Although ArtGS and REArtGS++ utilize two-stage observations, our method still achieve bests mean results on both CD-m and CD-s with 25.87, 31.02, significantly exceeding them. As shown in Fig. 9, our method yields high-fidelity part-level mesh reconstruction and accurate joint estimation. This demonstrates our part segmentation and joint estimation for planar Gaussians are simple but effective, achieving realistic results without two-stage observations.
To further evaluate the articulated modeling performance for multi-part objects, we conduct experiments on the ArtGS-Multi dataset. The quantitative and qualitative results are presented in Table. V and Fig. 11 respectively. Experimental results validate the robustness and superiority of our method for modeling multi-part articulated objects.
TABLE V: Quantitative results for articulated modeling on ArtGS-Multi dataset.
Metrics Method Storage Storage Table Table Oven Mean
47468 (7 parts) 45503 (4 parts) 25493 (4 parts) 31249 (5 parts) 101908 (4 parts)
Axis Ang. $\downarrow$ ArtGS [17] 65.40 82.10 45.20 38.70 50.30 56.34
REArtGS++ [32] 38.20 45.60 35.40 32.50 28.90 36.12
SPAGS (Ours) 4.20 6.80 3.50 4.10 2.90 4.30
Axis Pos. $\downarrow$ ArtGS [17] 2.05 2.30 - 1.15 1.70 1.80
REArtGS++ [32] 1.55 1.62 - 1.38 1.40 1.49
SPAGS (Ours) 0.14 0.22 - 0.11 0.08 0.14
CD-s $\downarrow$ ArtGS [17] 80.10 78.30 62.50 55.40 68.20 68.90
REArtGS++ [32] 52.40 48.50 45.20 42.60 46.80 47.10
SPAGS (Ours) 16.50 18.20 15.40 16.10 14.80 16.20
CD-m $\downarrow$ ArtGS [17] 195.80 212.40 185.60 178.50 188.20 192.10
REArtGS++ [32] 138.20 142.60 135.40 128.50 132.10 135.36
SPAGS (Ours) 22.40 26.50 20.80 19.50 18.60 21.56
CD $\downarrow$ ArtGS [17] 77.95 85.35 64.05 56.95 68.20 70.50
REArtGS++ [32] 30.30 50.55 45.30 40.55 24.45 38.23
SPAGS (Ours) 9.45 22.35 18.10 17.80 6.70 14.88
@@
TABLE VI: Ablation of view perception on PARIS dataset. Mean results are reported.
Settings CD $\downarrow$ CD-m $\downarrow$ CD-s $\downarrow$
@@
Optimal Perception 9.60 25.87 31.02
IV-C Ablation Study
Ablation on View Perception. We conduct the ablation study of view perception on PARIS dataset. Quantitative results are shown in Table. VI. “Random Sampling” denotes selecting 4-view images randomly from candidates and we take the average results over 10 trials. “Pre-defined” refers to manually choosing promising 4-view images covering 360 degrees around the objects.